a vollyball player hits a ball and it goes straight up into the air with an initial velocity of 16 feet per second from an initial height of 4 feet. find the function that describes the height of the ball in terms of the t and determine the time at which the ball reaches its maximum height.
Answer by stanbon(75887)

a vollyball player hits a ball and it goes straight up into the air with an initial velocity of 16 feet per second from an initial height of 4 feet. find the function that describes the height of the ball in terms of the t and determine the time at which the ball reaches its maximum height.
----
h(t) = -16t^2+16t + 4
----
Max height occurs when x = -b/(2a) = -16/(2*-16) = 1/2
Max height = h(1/2) = -16(1/4)+16(1/2)+4 = 8 ft
============
Cheers,
Stan H.